OpenStudy (anonymous):

square root of (x+6)^6

OpenStudy (tkhunny):

Whoops! Wolfram needs a better explanation. \(\sqrt{(x+6)^6} = \left|(x+3)^3\right|\) The left-hand expression is necessarily non-negative. The right-hand expression, without the Absolute Values, is not restricted from being negative. We must restrict it with the Absolute Values.